In Problems 23–32, assume that the random variable X is normally
distributed, with mean m = 50 and standard deviation s = 7.
Compute the following probabilities. Be sure to draw a normal
curve with the area corresponding to the probability shaded
24. P(X > 65)
28. P(56 < X < 68)
thank you very much!

Answer by ewatrrr(24785)

Population: m = 50 and standard deviation s = 7.
P(x > 65) = 1- P(z ≤ 15/7) = 1- P(z ≤ 2.1429) = 1 -.9839 = .016 0r 1.16%
P(x > 65): shade area to the right of z = 2.1429
for 28. P(z ≤ 2.5714) - P(z ≤ .8571) find z's and shade area between them
For the normal distribution: Below: z = 0, z = ± 1, z= ±2 , z= ±3 are plotted.
Area under the standard normal curve to the left of the particular z is P(z)
Note: z = 0 (x value: the mean) 50% of the area under the curve is to the left and 50% to the right